OpenStudy (anonymous):

You have on hand morphine 500 mg in 500 mL normal saline. You need to administer morphine 80 mg/hour in a continuous IV infusion to a patient. How many milliliters per hour will you set on the controller?

OpenStudy (anonymous):

You know that 500mg is related to 500mL; that is,\[500mg \iff 500mL\]But that's the same as (after dividing both sides by 500),\[1mg \iff 1mL\]but again, that's the same as\[\frac{1mg}{h} \iff \frac{1mL}{h}\]which again is the same as\[\frac{80mg}{h}\iff \frac{80mL}{h}\]

OpenStudy (anonymous):

You need to set the controller to 80mL/hour.

OpenStudy (anonymous):

how do i write this out to show the work?

OpenStudy (anonymous):

Well, pretty much as I did it.

OpenStudy (anonymous):

They're a set of equivalent statements.

OpenStudy (anonymous):

You could write it out word-for-word if you want to.